# GEM3D: GEnerative Medial Abstractions for 3D Shape Synthesis

### 26 Feb 2024 v1

@@ 4.4. Ablation studies
Figure 5. Category-conditioned shape generation on ShapeNet. We show generated shapes from  GEM3D for five categories: chair, lamp, airplane, table, bench and watercraft. Odd rows are skeletons generated by our model; even rows are surfaces sampled from them.
@@
Figure 8. Surface sampling conditioned on skeleton. Our model can sample diverse and structurally consistent surfaces from generated skeletons.
@@
Figure 9. Surface generation based on user-modeled skeletons. Our model is able to generalize to unseen structures encoded in user-provided skeletons, and produces plausible surface samples from them.
5. Conclusion
@@
Our method still has a number of limitations and offers avenues for further research. First, it relies on a simplified form of point-based skeletons with fixed resolution. Generalizing our method to create more expressive representations, such as skeletal diagrams (Guo et al., 2023), may further enhance the topology information encoded in our model. Despite the improvements in the topology of reconstructed shapes, our method does not guarantee certain topology characteristics; the surfaces might still mismatch a target topology (e.g., a certain genus) or contain undesired, noisy shape components. Finally, an interesting future avenue would be to enable interactive skeleton editing, exploration, and shape synthesis from the edits.
@@ Appendix A Per-category evaluation
F1 $\uparrow$ table 0.961 0.982 0.998 0.999 0.999 0.999 0.991
@@
Table 8. Evaluation of reconstructed surfaces on ShapeNet based on Chamfer Distance (CD), Intersection over Union (IoU), and F1 scores for the auto-encoding task. Here we use the original 3DShape2VecSet protocol. We report the same seven categories and numbers are not scaled as in their paper.

### 11 Apr 2024 v2

@@ 4.1. Category-driven shape generation
Figure 5. Category-conditioned shape generation on ShapeNet. We show generated shapes from  GEM3D for five categories: chair, lamp, airplane, table, bench and watercraft (from top to bottom). Odd rows are skeletons generated by our model; even rows are surfaces sampled from them.
Qualitative results
Figure 5 demonstrates generated samples from our method for various categories from ShapeNet. We observe that our method is able to generate structurally challenging patterns (e.g., grid-like patterns for chair backs), thin parts (e.g, wings for airplanes), tubular or generalized cylinder-like parts (e.g., lamp wires or other connecting pieces).
Timings
It takes around 40 seconds for GEM3D to generate the skeleton and surface, while 3DS2VS takes about 20 seconds. Reconstruction takes 6s for GEM3D vs 12s for 3DS2VS ($128^{3}$ grid; all measured on a NVidia RTX 2080Ti).
4.2. Surface reconstruction from point clouds
In this application, the input to our method is a sparse point cloud of a shape and the output is a reconstructed surface. Here, we also follow the “auto-encoding” setting also used in 3DILG (Zhang et al., 2022) and 3DS2VS (Zhang et al., 2023), where the point cloud is obtained after sampling the original input surface. In this autoencoding setting, we only use the second diffusion stage of our method i.e., given a point-based skeleton, the second stage generates the medial latents, then our surface decoder uses them to reconstruct the surface. To apply our method on this task, we need as input a point-sampled skeleton. To obtain such a skeleton, we use a variant of P2PNet (Yin et al., 2018), an encoder-decoder network that takes as input a point cloud of a shape’s surface and converts it to a point-based skeleton representation. Compared to the original P2PNet, we replaced the original PointNet++ encoder (Qi et al., 2017) with the cross-attention-based shape encoder from 3DS2VS (Zhang et al., 2023), including its positional-based embedding functions. We train this P2PNet variant in the training split of ShapeNet. Then at test time, we use its output skeletons as input to our model decoder. We emphasize that no manual input or supervision were needed to obtain the skeletons – P2PNet is trained in a unsupervised manner.
Baselines and metrics
We again compare our method with 3DILG (Zhang et al., 2022) and 3DS2VS (Zhang et al., 2023), which were also applied in the same auto-encoding setting. All methods use the same number of latent codes ($2048$).
For evaluation purposes, we compare reconstructed shapes to their corresponding reference (“ground-truth”) shapes, which the input point clouds were sampled from. Following 3DS2VS (Zhang et al., 2023), we use the Chamfer Distance (CD), Intersection Over Union (IoU) and F-Score (F1). Compared to 3DS2VS’ protocol, we made the following changes: (i) for computing IoU, we use a higher-resolution grid of $256^{3}$ instead of $128^{3}$ to better characterize reconstruction of small-scale surface details, thin parts, and holes, (ii) input point clouds are sampled with farthest point sampling instead to better cover small and thin parts. In the supplement, we provide comparisons using the original 3DS2VS’ protocol.
@@ 4.2. Surface reconstruction from point clouds
Metric Test dataset 3DILG 3DS2VS Ours
CD ($\downarrow$) Shapenet (10 cat.) 1.69 1.32 1.17
CD ($\downarrow$) Shapenet (all cat.) 2.05 1.70 1.38
IOU ($\uparrow$) Shapenet (10 cat.) 90.1 96.2 95.9
IOU ($\uparrow$) Shapenet (all cat.) 90.5 95.4 95.5
F1 ($\uparrow$) Shapenet (10 cat.) 97.7 98.5 99.3
F1 ($\uparrow$) Shapenet (all cat.) 96.3 97.4 98.7
Table 2. Point cloud surface reconstruction (“auto-encoding”) evaluation on ShapeNet. The rows “ShapeNet (10 cat.)” report the measures averaged over the $10$ largest categories of ShapeNet, while the rows “ShapeNet (all)” report the measures averaged over all $55$ ShapeNet categories. For evaluation per-category, please see the supplement.
Thingi10K #$\#shapes$ 3DILG 3DS2VS Ours
all 9997 2.12 1.74 1.68
$genus\geq 0$ 5608 1.93 1.62 1.61
$genus\geq 5$ 1037 2.41 1.96 1.78
$genus\geq 10$ 489 2.88 2.30 1.98
CD $\downarrow$ $genus\geq 20$ 229 3.67 2.91 2.32
all 9997 93.6 96.4 96.7
$genus\geq 0$ 5608 95.3 97.4 97.4
$genus\geq 5$ 1037 91.1 94.8 95.9
$genus\geq 10$ 489 87.5 92.3 93.7
F1 $\uparrow$ $genus\geq 20$ 229 82.6 88.0 90.4
Table 3. Evaluation measures on point cloud surface reconstruction in the Thingi10K dataset. We note that none of the methods are trained on Thingi10K. We report performance in terms of CD and F1 scores for shapes of increasing genus.
Experimental setup
Following (Zhang et al., 2022) and (Zhang et al., 2023), we use the same training and test split from ShapeNet-v2 to train and test our method respectively. We also test all competing methods in a more challenging out-of-distribution testing scenario: after training on ShapeNet, we test all methods on Thingi10K (Zhou and Jacobson, 2016). We note that the whole dataset is $10K$ test shapes, $4$ times larger than the ShapeNet’s test split. The Thingi10K contains man-made objects that often possess highly complex topology. Vast majority of the Thingi10K shapes do not exists in ShapeNet and do not even relate to any of its categories. Thingi10K also provides topological complexity information (genus) for a large subset of the dataset. We use it as a proxy of topological complexity for additional evaluation.
Quantitative results
Table 2 shows quantitative evaluation on ShapeNet-v2. The odd rows report the measures averaged over the $10$ largest categories of ShapeNet, and the even rows report the measures averaged over all $55$ categories. In terms of the surface-based metrics of CD and F1 scores, our model is consistently better than the baselines. In terms of IoU, our model has comparable performance. The results reflect a more accurate surface reconstruction, without compromising volumetric (IoU) accuracy. The supplement reports the performance for each of the largest $10$ categories – again, we outperform prior methods, especially for topologically challenging categories, such as lamps and watercrafts. The out-of-distribution evaluation on Thingi10K is shown in Table 3. Our model clearly outperforms the baselines – the performance gap increases for topologically challenging shapes with higher genus. We note that this evaluation does not include IoU since some Thingi10K ground-truth shapes are not watertight, thus, volume-based metrics cannot be accurately assessed.
@@
Figure 8. Surface sampling conditioned on skeleton. Our model can sample diverse and structurally consistent surfaces from generated skeletons. Surfaces were sampled using following category tokens (from top to bottom): chair, chair, lamp, lamp, airplane.
@@
Figure 9. Surface generation based on user-modeled skeletons. Our model is able to generalize to unseen structures encoded in user-provided skeletons, and produces plausible surface samples from them. Surfaces were sampled using following category tokens (from top to bottom): lamp, lamp, airplane, chair.
Qualitative results
Figures 6 and 7 shows qualitative evaluation on ShapeNet-v2 and Thingi10K respectively. In particular, the Thingi10K results show that our method can better reconstruct surface topology e.g., the connectivity of different shape parts and their surface holes. We suspect that this is due to the ability of our method to capture structure and topology information through the neural medial representation, which in turn guides the reconstruction. In contrast, 3DILG and 3D2VS tend to oversmooth surface details, miss connections, and close holes in the shapes.
Ablation studies
In our supplementary material, we also show the impact of various choices in our method, including testing with different number of medial points and using closest medial points versus closest envelopes for reconstruction.
4.3. Skeleton-driven shape synthesis
An alternative application of our method is to generate surfaces driven by an input skeleton in the form of a point-based MAT. From a practical point of view, one possibility for users of our method is to execute our first diffusion stage, obtain various shape structures represented in their MATs, then select the one matching their target structure or topology goals more. Then the user can execute the second stage to obtain various shape alternatives conforming to the given structure. Figure 8 highlights several examples of this application scenario. We show a generated MAT, then diverse surfaces conforming to it. For example, our method can generate diverse lamp stands from a table lamp structure or chairs of diverse part thicknesses conforming to a particular chair structure.
5. Discussion
@@
Figure 10. Failure cases. (Left) Generated MAT is unrealistic for lamp category. (Right) Generated surface is non-smooth.
Limitations
Failure cases of our method include unrealistic generated skeletons and noisy generated surfaces due to noise in MATs (Figure 10). We hypothesize this is due to either noisy training MATs or the training data being too limited for the model to learn comprehensive structural priors for challenging categories. Our method also has a number of other limitations. First, it relies on a simplified form of point-based skeletons. Second, the diversity of surfaces sampled from skeletons is sometimes limited. Investigation of less expressive skeletal representations or additional types of conditioning (e.g. images or text) can improve this aspect of our model. Other future directions would be to enable interactive skeleton editing, shape synthesis from edits, combinations of our model with sketch-based editing models (Luo et al., 2023) and physics-based shape deformations (Lan et al., 2020).
@@ Appendix A Per-category evaluation
mean (selected) 0.822 0.881 0.929 0.952 0.957 0.967 0.954
mean (all) 0.825 0.888 0.934 0.953 0.955 0.965 0.944
table 0.961 0.982 0.998 0.999 0.999 0.999 0.991
@@
mean (selected) 0.898 0.951 0.975 0.976 0.977 0.979 0.988
F1 $\uparrow$ mean (all) 0.858 0.933 0.967 0.966 0.966 0.970 0.967
Table 8. Evaluation of reconstructed surfaces on ShapeNet based on Chamfer Distance (CD), Intersection over Union (IoU), and F1 scores for the auto-encoding task. Here we use the original 3DShape2VecSet protocol. We report the same seven categories and averages (numbers are not scaled as in their paper).
@@ Appendix C Skeletonization
Appendix D Model capacity
Comparison between GEM3D and 3DS2VS in terms of number of parameters is shown in Table 10. Despite having two diffusion stages, total number of parameters for GEM3D model is only 1.26x more to those of 3DS2VS. This is due to fact that our model has way simpler surface decoder. We didn’t have to run hyperparameter optimization for our generative models but we hypothesize that it can be potentially substantially reduced, especially for surface decoder.
Decoder Gen-MAT Gen-Surf Total
3DS2VS 102M - 164.2M 266.2M
GEM3D 6.3M 164.9M 164.9M 336.1M
Table 10. Parameter comparison between 3D2SVS vs GEM3D. Despite having two diffusion stages, overall increase in parameters for our model is only 1.26x compared to 3DS2VS.